# Instance-Wise Adaptive Sampling for Dataset Construction in Approximating Inverse Problem Solutions

### v1

@@ Abstract
Keywords: inverse problems, adaptive sampling, scientific machine learning, data efficiency
@@ 1 Introduction
Inverse problems represent a fundamental class of challenges across numerous scientific and engineering domains, where the goal is to infer underlying parameters or structures from observable measurements. These problems are often notoriously difficult due to their ill-posed nature, often requiring sophisticated mathematical techniques and substantial computational resources to solve effectively. In recent years, deep learning approaches have emerged as powerful tools for approximating solutions to inverse problems, offering the potential for significantly faster inference while achieving reasonable accuracy compared to traditional optimization-based methods; see, e.g., the recent reviews (Arridge et al., 2019; Ongie et al., 2020; Ying, 2022) and references therein.
However, a critical limitation of deep learning approaches for inverse problems is their considerable data hunger. Training effective inverse maps from measurements to underlying parameters based on neural networks typically requires large datasets of input-output pairs (Zhou et al., 2023; Klug and Heckel, 2023; Adcock et al., 2024), which can be prohibitively expensive to collect and use, for instance, when each forward simulation involves solving complex partial differential equations (PDEs). As prior knowledge about the inferred parameters becomes less constrained, the data requirements become increasingly demanding because more data is necessary to sufficiently cover the parameter space. This creates a substantial obstacle for applying deep learning to realistic inverse problems with complex, high-dimensional parameter spaces.
@@
We demonstrate our method on an inverse scattering problem for the Helmholtz equation (Colton et al., 1998; Kirsch, 2011), a challenging inverse problem with applications in radar, sonar, medical imaging, and seismic exploration. In this context, the goal is to determine the properties of an unknown heterogeneous medium by probing it with incident waves and measuring the resulting scattered waves at distant locations. Numerical experiments show that models trained with our adaptive sampling approach can achieve performance comparable to or better than models trained on datasets many times larger. For a single challenging instance, the required sample size can be reduced by one to two orders of magnitude, depending on the complexity of the parameters to infer.
The proposed instance-wise adaptive sampling strategy can also be viewed as a form of inference-time scaling in inverse problems, where computational resources for data generation are allocated more efficiently by focusing on the most relevant regions of the parameter space during inference time. A similar shift is emerging in large language models (LLMs), where further scaling of pre-training is increasingly constrained by the scarcity of data and computational resources (Villalobos et al., 2024; Muennighoff et al., 2025). As a result, there is growing interest in methods that adapt model behavior or resource allocation at inference time, on a per-query basis (Snell et al., 2024; Liu et al., 2025; OpenAI, 2024).
@@ 2 Methodology
Although in many setups the parameter $q$ lives in a high-dimensional ambient space, e.g. $\mathbb{R}^{N_{2}}$ for some large $N_{2}$, it is often the case that we have prior knowledge that $q$ lies on or close to some potentially low-dimensional manifold $\mathcal{M}$ in $\mathbb{R}^{N_{2}}$. In particular, the intrinsic dimension $N_{1}$ of the parameter manifold $\mathcal{M}$ may be much smaller than the dimension $N_{2}$ of the ambient space for some applications. Such prior knowledge could come either from the underlying physics or from the fact that the inverse problem is so ill-conditioned that only limited information about $q$ can be reliably reconstructed (Bal and Ren, 2009). In this paper, we consider two representative classes of priors: smoothness-based and geometry-based, both of which are described in detail in Section 3.1.
@@ 2.1 Analogy with Inference-Time Compute
The proposed method is part of a broader trend in machine learning that shifts more computation to the inference stage, a direction that has gained significant traction in the context of large language models (LLMs) (Snell et al., 2024; Liu et al., 2025; OpenAI, 2024). For LLMs, inference-time computation typically falls into two main categories, as illustrated in (Snell et al., 2024, Figure 5): (1) parallel sampling (Brown et al., 2024; Stroebl et al., 2024), and (2) sequential revision (Madaan et al., 2024; Qu et al., 2025; Welleck et al., 2022), with recent work also exploring hybrids of both approaches.
@@
On the other hand, in sequential revision, the LLM first generates an initial solution and then iteratively refines its answer. While our instance-wise adaptive sampling method does not precisely align with existing LLM inference-time paradigms, it shares structural similarities with the sequential revision framework. In what follows, we draw a concrete analogy using Self-Refine approach introduced in Madaan et al. (2024) as a representative example.
For ease of explanation, we reproduce the pseudocode of Self-Refine from (Madaan et al., 2024, Algorithm 1) as Algorithm 2. The method begins with a preliminary generation of the answer from the LLM, followed by a feedback step in which the same model critiques the answer, and a refinement step in which the model incorporates the feedback to produce an improved version. This process is repeated for multiple rounds. Few-shot examples are used in the prompt to guide the model during generation, feedback, and refinement, denoted in Algorithm 2 by $p_{\text{gen}}$, $p_{\text{fb}}$, and $p_{\text{refine}}$ respectively, with $||$ indicating prompt concatenation.
Algorithm 2 LLM Self-Refine (Madaan et al., 2024, Algorithm 1)
@@
Table 1: Comparison between LLM Self-Refine (Madaan et al., 2024) and our adaptive sampling method for inverse problems.
@@ 2.2 Other Related Works
There has also been related work in the applied mathematics and scientific computing communities. Perhaps the closest to our setting is Tatsuoka et al. (2025), which introduces an instance-wise adaptive refinement method in the context of Bayesian inverse problems. Their objective is to characterize the full posterior distribution of the parameter, which leads them to focus on low-dimensional parameter spaces (one- or two-dimensional). Their method also involves only two sampling levels, whereas ours allows multiple rounds of refinement for high-dimensional parameters. Overall, the two approaches share some spirit at a high level but are not directly comparable.
The remaining related work can be broadly divided into two main categories. One line studies other machine learning approaches to inverse problems. For example, Melia et al. (2025) proposes a neural network architecture for the multi-frequency inverse scattering problem, first constructing a low-frequency approximation from the lowest-frequency measurement and then iteratively refining it with higher-frequency data. Another recent work Jiang et al. (2024) uses reinforcement learning to adaptively choose sensor locations and incident frequencies, highlighting the benefits of adaptivity in frequency design for inverse scattering. By contrast, we focus on the single-frequency case, and our progressive refinement is not frequency-based. Despite the difference, the analogy in their refinement structure suggests an interesting future direction: frequency-based adaptive sampling, where early rounds reconstruct low-frequency components that are subsequently refined into higher-frequency approximations.
The second category applies sequential adaptive sampling to settings outside inverse problems, such as adaptive collocation point selection for physics-informed neural networks (Lu et al., 2021; Wu et al., 2023) and adaptive proposal construction for rare event probability estimation (Tong and Stadler, 2023). Although the applications and objectives differ, these methods share a structural similarity with ours: at each round, random samples are drawn adaptively based on the current state, used to update the state, and repeated over multiple rounds. The successes in these areas highlight the versatility of sequential adaptive sampling and suggest that it could be fruitfully explored in still more domains.
@@ 3 Example Problem: Inverse Scattering
where $m$ is the given measurement data. It is important to note that the number of basis functions in $\mathcal{Q}_{N_{3}}$ may be significantly smaller than that needed to fully resolve the fine-scale features of the ground-truth field $q^{*}$. This deliberate restriction of the search space helps mitigate the ill-posedness inherent in the inverse scattering problem, which is fundamentally constrained by the Heisenberg uncertainty principle (Chen, 1997). A similar strategy was adopted in Borges et al. (2017); Askham and Borges (2024) to address the same challenge. Note that while the restricted search space $\mathcal{Q}_{N_{3}}$ can be seen as prior information on $q^{*}$, in our setup, the prior manifold $\mathcal{M}$ is a separate space and is not necessarily related to the smoothness of $q^{*}$. Two examples of the prior manifold $\mathcal{M}$ are discussed in the next subsection.
@@ 3.1 Prior Knowledge of the Parameter
In this setting, the projection onto $\mathcal{M}$ in line 4 and the local perturbation on $\mathcal{M}$ in line 6 of Algorithm 1 can be implemented as follows. Given the prediction $\widehat{q}^{(t)}$ of the current neural network model $\mathcal{NN}_{\theta_{t}}$, we utilize the phase-coding method (Atherton and Kerbyson, 1999), implemented in the imfindcircles function in MATLAB, to detect all the possible disks $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$ in $\widehat{q}^{(t)}$. The field $\widehat{q}^{(t)}$ is then averaged over each disk $\mathcal{D}_{i}$ to obtain an associated amplitude $a_{i}$. The collection of disks $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$ together with their amplitudes $a_{1},\dots,a_{n}$ corresponds to the projection onto the prior manifold in line 4 of Algorithm 1. We then sample around $\widehat{q}^{(t)}$ on the parameter manifold $\mathcal{M}$ by randomly perturbing the centers, radii, and amplitudes of each of the disks.
@@ 4 Numerical Results
We choose to implement the construction of the adaptive datasets in line 9 of Algorithm 1 by combining the adaptively sampled local dataset of size $N_{\text{adapt}}$ with the $N_{\text{base}}$ elements in the base model dataset whose parameter fields are closest to the current prediction, measured in terms of the $\ell^{2}$ norm in $\mathbb{R}^{N_{3}^{2}}$. We note that this way of combining local data and base data, along with the specific choices of $N_{\text{adapt}}$ and $N_{\text{base}}$ in Tables 3 and 4, are current design choices; alternative configurations can certainly be explored.
@@
To systematically quantify the improvement of the adaptive method over the non-adaptive method, we run the adaptive method for a fixed number of rounds $N_{\text{round}}$ on all test data, rather than using a data-dependent stopping criterion as in Algorithm 1. This allows us to track the average relative error $\varepsilon_{\text{rel}}$ after each round. The value of $N_{\text{round}}$ is chosen so that the average error across all test data plateaus, as illustrated in the left panels of Figures 5 and 8. Note that in practice, we do not have access to the reconstruction error because we do not know the ground-truth field $q^{*}$. Instead, we can track the measurement error, i.e., the relative $\ell^{2}$ error in $\mathbb{R}^{N_{d}\times N_{t}}$ between the measurement associated to the reconstructed parameter $\widehat{q}$ and the given measurement. A practical stopping criterion is to terminate when the measurement error begins to plateau.
@@ 5 Discussion and Future Work
While the proposed adaptive sampling framework demonstrates strong performance in solving inverse scattering problems under structured priors, several important directions remain for future exploration. First, our current numerical experiments assume noiseless measurement data. Investigating robustness under various noise levels is, therefore, a natural next step. Second, the adaptive sampling strategy for learning the inverse map is not limited to the inverse scattering setup considered here; it can be readily applied to other inverse problems, such as wave inversion (Wu and Lin, 2019; Ding et al., 2025), or combined with classical approaches like the direct sampling method (Ning et al., 2023, 2025). Finally, we have focused on priors defined by a manifold $\mathcal{M}$, relying on explicit structural assumptions. However, in many practical scenarios, prior information is more realistically described by a distribution or density supported on $\mathcal{M}$, which can be learned from data. Modeling such distributions based on available datasets allows us to go beyond rigid manifold assumptions, providing richer and more flexible prior information that may improve both sample efficiency and generalization. In this context, generative modeling techniques such as score-based diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) offer a promising direction for representing and sampling from complex prior or posterior distributions (Chung et al., 2023; Bruna and Han, 2024; Zhang et al., 2025). Exploring these extensions may further enhance the practicality and expressiveness of the adaptive sampling approach.
Acknowledgments and Disclosure of Funding

### v2

@@ Abstract
Key words. inverse problems, adaptive sampling, scientific machine learning, data efficiency
@@ 1 Introduction
Inverse problems represent a fundamental class of challenges across numerous scientific and engineering domains, where the goal is to infer underlying parameters or structures from observable measurements. These problems are often notoriously difficult due to their ill-posed nature, often requiring sophisticated mathematical techniques and substantial computational resources to solve effectively. Traditional optimization-based methods for challenging inverse problems are often highly sensitive to initialization and can converge to undesirable local minimizers when the initial guess is insufficiently accurate [34, 9, 41, 21]. In recent years, deep learning approaches have emerged as powerful tools for approximating solutions to inverse problems, offering the potential for significantly faster inference while achieving reasonable accuracy compared to traditional optimization-based methods; see, e.g., the recent reviews [2, 25, 39] and references therein.
However, a critical limitation of deep learning approaches for inverse problems is their considerable data hunger. Training effective inverse maps from measurements to underlying parameters based on neural networks typically requires large datasets of input-output pairs [41, 17, 1], which can be prohibitively expensive to collect and use, for instance, when each forward simulation involves solving complex partial differential equations (PDEs). As prior knowledge about the inferred parameters becomes less constrained, the data requirements become increasingly demanding because more data is necessary to sufficiently cover the parameter space. This creates a substantial obstacle for applying deep learning to realistic inverse problems with complex, high-dimensional parameter spaces.
@@
We demonstrate our method on an inverse scattering problem for the Helmholtz equation [12, 16], a challenging inverse problem with applications in radar, sonar, medical imaging, and seismic exploration. In this context, the goal is to determine the properties of an unknown heterogeneous medium by probing it with incident waves and measuring the resulting scattered waves at distant locations. Numerical experiments show that models trained with our adaptive sampling approach can achieve performance comparable to or better than models trained on datasets many times larger. For a single challenging instance, the required sample size can be reduced by one to two orders of magnitude, depending on the complexity of the parameters to infer.
The proposed instance-wise adaptive sampling strategy can also be viewed as a form of inference-time scaling in inverse problems, where computational resources for data generation are allocated more efficiently by focusing on the most relevant regions of the parameter space during inference time. A similar shift is emerging in large language models (LLMs), where further scaling of pre-training is increasingly constrained by the scarcity of data and computational resources [35, 22]. As a result, there is growing interest in methods that adapt model behavior or resource allocation at inference time, on a per-query basis [28, 18, 26].
@@ 2 Methodology
Although in many setups the parameter $q$ lives in a high-dimensional ambient space, e.g. $\mathbb{R}^{N_{2}}$ for some large $N_{2}$, it is often the case that we have prior knowledge that $q$ lies on or close to some potentially low-dimensional manifold $\mathcal{M}$ in $\mathbb{R}^{N_{2}}$. In particular, the intrinsic dimension $N_{1}$ of the parameter manifold $\mathcal{M}$ may be much smaller than the dimension $N_{2}$ of the ambient space for some applications. Such prior knowledge could come either from the underlying physics or from the fact that the inverse problem is so ill-conditioned that only limited information about $q$ can be reliably reconstructed [5]. In this paper, we consider two representative classes of priors: smoothness-based and geometry-based, both of which are described in detail in Section 3.1.
@@ 2.1 Analogy with Inference-Time Compute
The proposed method is part of a broader trend in machine learning that shifts more computation to the inference stage, a direction that has gained significant traction in the context of large language models (LLMs) [28, 18, 26]. For LLMs, inference-time computation typically falls into two main categories, as illustrated in [28, Figure 5]: (1) parallel sampling [7, 31], and (2) sequential revision [20, 27, 36], with recent work also exploring hybrids of both approaches.
@@
On the other hand, in sequential revision, the LLM first generates an initial solution and then iteratively refines its answer. While our instance-wise adaptive sampling method does not precisely align with existing LLM inference-time paradigms, it shares structural similarities with the sequential revision framework. In what follows, we draw a concrete analogy using Self-Refine approach introduced in [20] as a representative example.
For ease of explanation, we reproduce the pseudocode of Self-Refine from [20, Algorithm 1] as Algorithm 2. The method begins with a preliminary generation of the answer from the LLM, followed by a feedback step in which the same model critiques the answer, and a refinement step in which the model incorporates the feedback to produce an improved version. This process is repeated for multiple rounds. Few-shot examples are used in the prompt to guide the model during generation, feedback, and refinement, denoted in Algorithm 2 by $p_{\text{gen}}$, $p_{\text{fb}}$, and $p_{\text{refine}}$ respectively, with $||$ indicating prompt concatenation.
Algorithm 2 LLM Self-Refine [20, Algorithm 1]
@@
Table 1: Comparison between LLM Self-Refine [20] and our adaptive sampling method for inverse problems.
@@ 2.2 Other Related Works
There has also been related work in the applied mathematics and scientific computing communities. Perhaps the closest to our setting is [32], which introduces an instance-wise adaptive refinement method in the context of Bayesian inverse problems. Their objective is to characterize the full posterior distribution of the parameter, which leads them to focus on low-dimensional parameter spaces (one- or two-dimensional). Their method also involves only two sampling levels, whereas ours allows multiple rounds of refinement for high-dimensional parameters. Overall, the two approaches share some spirit at a high level but are not directly comparable.
The remaining related work can be broadly divided into two main categories. One line studies other machine learning approaches to inverse problems. For example, [21] proposes a neural network architecture for the multi-frequency inverse scattering problem, first constructing a low-frequency approximation from the lowest-frequency measurement and then iteratively refining it with higher-frequency data. Another recent work [15] uses reinforcement learning to adaptively choose sensor locations and incident frequencies, highlighting the benefits of adaptivity in frequency design for inverse scattering. By contrast, we focus on the single-frequency case, and our progressive refinement is not frequency-based. Despite the difference, the analogy in their refinement structure suggests an interesting future direction: frequency-based adaptive sampling, where early rounds reconstruct low-frequency components that are subsequently refined into higher-frequency approximations.
The second category applies sequential adaptive sampling to settings outside inverse problems, such as adaptive collocation point selection for physics-informed neural networks [19, 37] and adaptive proposal construction for rare event probability estimation [33]. Although the applications and objectives differ, these methods share a structural similarity with ours: at each round, random samples are drawn adaptively based on the current state, used to update the state, and repeated over multiple rounds. The successes in these areas highlight the versatility of sequential adaptive sampling and suggest that it could be fruitfully explored in still more domains.
@@ 3 Example Problem: Inverse Scattering
where $m$ is the given measurement data. It is important to note that the number of basis functions in $\mathcal{Q}_{N_{3}}$ may be significantly smaller than that needed to fully resolve the fine-scale features of the ground-truth field $q^{*}$. This deliberate restriction of the search space helps mitigate the ill-posedness inherent in the inverse scattering problem, which is fundamentally constrained by the Heisenberg uncertainty principle [10]. A similar strategy was adopted in [6, 3] to address the same challenge. Note that while the restricted search space $\mathcal{Q}_{N_{3}}$ can be seen as prior information on $q^{*}$, in our setup, the prior manifold $\mathcal{M}$ is a separate space and is not necessarily related to the smoothness of $q^{*}$. Two examples of the prior manifold $\mathcal{M}$ are discussed in the next subsection.
@@ 3.1.1 Disk prior
In this setting, the projection onto $\mathcal{M}$ in line 4 and the local perturbation on $\mathcal{M}$ in line 6 of Algorithm 1 can be implemented as follows. Given the prediction $\widehat{q}^{(t)}$ of the current neural network model $\mathcal{NN}_{\theta_{t}}$, we utilize the phase-coding method [4], implemented in the imfindcircles function in MATLAB, to detect all the possible disks $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$ in $\widehat{q}^{(t)}$. The field $\widehat{q}^{(t)}$ is then averaged over each disk $\mathcal{D}_{i}$ to obtain an associated amplitude $a_{i}$. The collection of disks $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$ together with their amplitudes $a_{1},\dots,a_{n}$ corresponds to the projection onto the prior manifold in line 4 of Algorithm 1. We then sample around $\widehat{q}^{(t)}$ on the parameter manifold $\mathcal{M}$ by randomly perturbing the centers, radii, and amplitudes of each of the disks.
@@ 4 Numerical Results
We choose to implement the construction of the adaptive datasets in line 9 of Algorithm 1 by combining the adaptively sampled local dataset of size $N_{\text{adapt}}$ with the $N_{\text{base}}$ elements in the base model dataset whose parameter fields are closest to the current prediction, measured in terms of the $\ell^{2}$ norm in $\mathbb{R}^{N_{3}^{2}}$. We note that this way of combining local data and base data, along with the specific choices of $N_{\text{adapt}}$ and $N_{\text{base}}$ in Tables 3 and 5, are current design choices; alternative configurations can certainly be explored.
@@
To systematically quantify the improvement brought by adaptive sampling for data construction within the learning-based method over its non-adaptive counterpart, as well as over traditional optimization-based approaches, we run the adaptive method for a fixed number of rounds $N_{\text{round}}$ on all test data, rather than using a data-dependent stopping criterion as in Algorithm 1. This allows us to track the average relative error $\varepsilon_{\text{rel}}$ after each round. The value of $N_{\text{round}}$ is chosen so that the average error across all test data plateaus, as illustrated in the left panels of Figures 5 and 8. Note that in practice, we do not have access to the reconstruction error because we do not know the ground-truth field $q^{*}$. Instead, we can track the measurement error, i.e., the relative $\ell^{2}$ error in $\mathbb{R}^{N_{d}\times N_{t}}$ between the measurement associated to the reconstructed parameter $\widehat{q}$ and the given measurement. A practical stopping criterion is to terminate when the measurement error begins to plateau.
@@ 4.1 Disk prior
We further compare the adaptive sampling approach with classical optimization-based methods; the results are summarized in Table 4. Here, the adaptive model is trained on the adaptively constructed dataset, while the base model is trained on the initial dataset. The optimization problem (4) is solved using Gauss–Newton, following [6]. Since Gauss–Newton is sensitive to initialization, we report results with two strategies: nearest-neighbor initialization and base-model initialization. For nearest-neighbor initialization, we select the closest measurement in the base training set and use its associated parameter field as the initial guess. As shown in Table 4, direct prediction from the adaptive model significantly reduces reconstruction error and outperforms Gauss–Newton even when initialized with the base-model prediction, indicating that the gain comes from adaptive data construction rather than post hoc optimization.
$N_{\text{disk}}\in[1,3]$ $N_{\text{disk}}\in[4,6]$
Direct prediction (base model) 30.2% 33.3%
Direct prediction (adaptive model) 9.3% 12.0%
Gauss–Newton (nearest-neighbor init.) 29.4% 66.3%
Gauss–Newton (base-model init.) 26.1% 29.1%
Table 4: Average reconstruction error on the test set under the disk prior, for direct prediction using the base and adaptive model (trained on the adaptively constructed dataset), and for Gauss–Newton refinement under two initialization strategies. Direct prediction from the adaptive model substantially outperforms both the base model and Gauss–Newton, even when the latter is initialized with the base-model prediction.
@@ 4.2 Fourier Prior
Additionally, for the Fourier prior, we also compare the adaptive sampling approach with classical optimization-based methods; the results are shown in Table 6. The same trend observed under the disk prior persists: direct prediction from the adaptive model yields substantially lower reconstruction error and outperforms Gauss–Newton, even when initialized with the base-model prediction.
$N_{F}=3$ $N_{F}=4$
Direct prediction (base model) 44.4% 53.6%
Direct prediction (adaptive model) 18.7% 35.7%
Gauss–Newton (nearest-neighbor init.) 97.6% 102.0%
Gauss–Newton (base-model init.) 42.9% 47.8%
Table 6: Average reconstruction error on the test set under the Fourier prior, for direct prediction using the base and adaptive model (the latter trained on the adaptively constructed dataset), and for Gauss–Newton refinement under two initialization strategies. As in the disk prior case, direct prediction from the adaptive model outperforms both the base model and Gauss–Newton, even when the latter is initialized with the base-model prediction.
@@ 5 Discussion and Future Work
While the proposed adaptive sampling framework demonstrates strong performance in solving inverse scattering problems under structured priors, several important directions remain for future exploration. First, our current numerical experiments assume noiseless measurement data. Investigating robustness under various noise levels is, therefore, a natural next step. Second, the adaptive sampling strategy for learning the inverse map is not limited to the inverse scattering setup considered here; it can be readily applied to other inverse problems, such as wave inversion [38, 13], or combined with classical approaches like the direct sampling method [23, 24]. Finally, we have focused on priors defined by a manifold $\mathcal{M}$, relying on explicit structural assumptions. However, in many practical scenarios, prior information is more realistically described by a distribution or density supported on $\mathcal{M}$, which can be learned from data. Modeling such distributions based on available datasets allows us to go beyond rigid manifold assumptions, providing richer and more flexible prior information that may improve both sample efficiency and generalization. In this context, generative modeling techniques such as score-based diffusion models [29, 14, 30] offer a promising direction for representing and sampling from complex prior or posterior distributions [11, 8, 40]. Exploring these extensions may further enhance the practicality and expressiveness of the adaptive sampling approach.
Acknowledgment